You have prepared a drug using a 4 g in 250 mL 0.9% NS solution for your patient with an irregular heart rhythm. The physician has ordered 6mg/min IV in a continuous infusion via an infusion pump. Calculate the mL/hour flow rate. Use dimensional analysis.
Answer #1

ANSWER.

* Calculation of flow rate.

        Flow rate = Volume(ml) / Time(hr)

Calculation of volume(ml)

         Volume(ml) = Desired dose(D) / Dose in hand (H) * Quantity.

Here

        Desired dose(D) = 6 mg/min

        Dose in hand (H) = 4 g.

   Convert gm to mg for calculation

                  1 gm = 1000 mg.

     Dose in hand (H) = 4000 mg.

            Quantity (Q) = 250 ml

As per above formula

        Volume(ml) = 6/4000 * 250

                            = 0.0015 * 250

                            = 0.38

                            = 0.4 ml/min

     Convert ml/min to ml/hr by multiply with 60

          So

              Flow rate(ml/hr) = 0.4 * 60

                                         = 24 ml/hr.